# Improving Pronunciation and Accent Conversion through Knowledge Distillation And Synthetic Ground-Truth from Native TTS

### v1

@@ I Introduction
Conventional AC [1, 2] methods require reference utterances in the target accent during synthesis, limiting their applicability due to the challenge of obtaining linguistically identical references in different accents. This research, like recent studies [3, 4], focuses on reference-free AC, which converts accents without needing reference utterances during inference. Reference-free AC methods generally fall into two categories: autoregressive and non-autoregressive architectures.
Autoregressive AC models , typically based on sequence-to-sequence (seq2seq) architectures, rely mostly on parallel data for training. However, such data—utterances from the same speakers in different accents—is scarce, making it difficult to obtain. To address this, data augmentation techniques like text-to-speech (TTS) [5] or voice conversion [6] are used to generate ground-truth audio for non-native speakers. These ground-truth audios, originating from different speakers, often differ in duration and prosody compared to the original non-native audio. While the encoder-decoder model with attention mechanisms can help aligning inputs and outputs, the differences in length make these models less suitable for applications requiring precise time synchronization, such as video dubbing. Additionally, there are some autoregressive models[7, 8] that trained with non-parallel data, but still share similar drawbacks, including slow inference speeds and unstable attention mechanism.
@@
Overall, to further enhance the AC model’s training process, we introduced several key improvements. In the first stage, we simultaneously pre-train the AC model while training the VITS-based native TTS, allowing the AC model to learn the distribution of native audio data comprehensively. After training the native VITS-based TTS, we leverage it to generate ideal ground-truth data for non-native audio, as described in the next section. In the second stage, we fine-tune the pre-trained AC model using the non-native input alongside the generated ground-truth output and also distillate knowledge from the native TTS. In summary, we train the VITS-based TTS model to assist in initializing the weights, generating ideal ground-truth data, and facilitating knowledge distillation for the AC model.
@@ IV Conclusion
In this work, our work presents a significant advancement in accent conversion and native-like pronunciation correction. Through the introduction of a two-stage training process that leverages VITS framework, we enable the AC model to better capture the accent-independent feature of speech while maintaining the speaker’s identity. The use of generated ground-truth data, alongside knowledge distillation from native TTS, further enhances the system’s ability to correct non-native pronunciations effectively. Objective and subjective evaluations demonstrate that our approach not only improves accent conversion but also addresses pronunciation issues, providing a more natural and comprehensible output. Future work may explore further refinement in emotion, prosody and accent feature to build upon these promising results.

### v2

@@ I Introduction
Conventional AC [1, 2] methods require reference utterances in the target accent during synthesis, limiting their applicability due to the challenge of obtaining linguistically identical references in different accents. This research, like recent studies [3, 4, 5, 6, 7], focuses on reference-free AC, which converts accents without needing reference utterances during inference. Reference-free AC methods generally fall into two categories: autoregressive and non-autoregressive architectures.
Autoregressive AC models , typically based on sequence-to-sequence (seq2seq) architectures, rely mostly on parallel data for training. However, such data—utterances from the same speakers in different accents—is scarce, making it difficult to obtain. To address this, data augmentation techniques like text-to-speech (TTS) [8] or voice conversion [9] are used to generate ground-truth audio for non-native speakers. These ground-truth audios, originating from different speakers, often differ in duration and prosody compared to the original non-native audio. While the encoder-decoder model with attention mechanisms can help aligning inputs and outputs, the differences in length make these models less suitable for applications requiring precise time synchronization, such as video conference dubbing [10, 11, 12]. Additionally, there are some autoregressive models[13, 14] that trained with non-parallel data, but still share similar drawbacks, including slow inference speeds and unstable attention mechanism.
@@
Overall, to further enhance the AC model’s training process, we introduced several key improvements. In the first stage, we simultaneously pre-train the AC model while training the VITS-based native TTS, allowing the AC model to learn the distribution of native audio data comprehensively. After training the native VITS-based TTS, we leverage it to generate ideal ground-truth data for non-native audio, as described in the next section. In the second stage, we fine-tune the pre-trained AC model using the non-native input alongside the generated ground-truth output and also distillate knowledge from the native TTS. In summary, we train the VITS-based TTS model to assist in initializing the weights, generating ideal ground-truth data, and facilitating knowledge distillation for the AC model.”originatingScript”:”m2”,”payload”:”guid”:”9a5f9d28-6c94-4bff-a24e-b5e10ecf90e15b8e9d”,”muid”:”1a8ee153-6a60-4b16-82ea-42e185928d0a78c2dd”,”sid”:”ea8eb107-bf17-4e70-b563-aa64cdb85f85496a83”
@@ IV Conclusion
Our work presents a significant advancement in accent conversion and native-like pronunciation correction. Through the introduction of a two-stage training process that leverages VITS framework, we enable the AC model to better capture the accent-independent feature of speech while maintaining the speaker’s identity. The use of generated ground-truth data, alongside knowledge distillation from native TTS, further enhances the system’s ability to correct non-native pronunciations effectively. Objective and subjective evaluations demonstrate that our approach not only improves accent conversion but also addresses pronunciation issues, providing a more natural and comprehensible output.
V Acknowledgment
This research was supported by a grant from Zoom Video Communications , Inc. T; European Commission Project Meetween (101135798), the Federal Ministry of Education and Research (BMBF) of Germany under the number 01EF1803B (RELATER), and the pilot program Core-Informatics of the Helmholtz Association (HGF). The work is enabled by the Horeka supercomputer (BMBF) and the DeltaGPU at NCSA/ACCESS grant[31].